OpenStudy (anonymous):

sum of 2 consecutive even numbers is 2096 what are the two numbers? show equation and how to solve it please, thank you!

OpenStudy (amistre64):

2096/2 is a good start

OpenStudy (anonymous):

no i need it like for example x(x+2)=2096 or something like that

OpenStudy (anonymous):

hello? you there?

OpenStudy (asnaseer):

are you sure its the sum of two and not the product of two consecutive even numbers?

OpenStudy (anonymous):

Yes the product omg thank you for correcting me!!

OpenStudy (asnaseer):

ok, so if the first number is even, then lets call it \(2x\). then the next consecutive even number would be \(2x+2\). so you need to solve:\[2x(2x+2)=2096\]

OpenStudy (anonymous):

oh ok thank you but what if it was the product of 2 consecutive odd numbers?

OpenStudy (asnaseer):

odd numbers can be written as \(2x+1\)

OpenStudy (anonymous):

isnt it the same which is 2x(x+2)?

OpenStudy (asnaseer):

\(2x\) always gives you an even number for x=1,2,3,... \(2x+1\) always gives you and odd number for x=1,2,3,...

OpenStudy (amistre64):

2x - 1 as well if thats easier

OpenStudy (anonymous):

oh ok thanks, my teacher said it would be the same as an odd though?

OpenStudy (asnaseer):

so, if you are told that the product of two consecutive odd numbers is 2096, then you could write:\[(2x+1)(2x+3)=2096\]

OpenStudy (anonymous):

oh ok thanks

OpenStudy (asnaseer):

amistre is right, \(2x-1\) is probably a better one to use.

OpenStudy (anonymous):

what about even?

OpenStudy (asnaseer):

are we going in circles here?

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

even is 2x(2x+2)?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

okay thanks let me write that down as a note. so 2 even consecutives is 2096 would be written as 2x(2x+2) and 2 consecutive odd would be (2x+1)(2x+3)=2096???.

OpenStudy (asnaseer):

yes, I would probably amistre's suggestion for odd numbers and use:\[(2x-1)(2x+1)=2096\]

OpenStudy (asnaseer):

*probably use

OpenStudy (anonymous):

oh iight. okay thanks!!

OpenStudy (asnaseer):

yw